OpenStudy (anonymous):

The graph of f(x) = 2^x + 4 shifts six units to the right when it is replaced with the graph of f(x) = 2^x - k. What is the value of k?

OpenStudy (anonymous):

2 4 6 10

OpenStudy (anonymous):

@jim_thompson5910 do you think you could help?

jimthompson5910 (jim_thompson5910):

the x+4 is all in the exponent?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

so is x-k ?

OpenStudy (anonymous):

yep

jimthompson5910 (jim_thompson5910):

to shift f(x) 6 units to the left, we replace every x with x+6

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be x+6+4?

jimthompson5910 (jim_thompson5910):

Correct \[\Large f(x) = 2^{x+4}\] \[\Large f(x+6) = 2^{x+6+4}\]

jimthompson5910 (jim_thompson5910):

simplify x+6+4 and compare it to x-k to figure out what k would be

OpenStudy (anonymous):

so it'll be x+10

OpenStudy (anonymous):

so k=10?

jimthompson5910 (jim_thompson5910):

x+10 x-k k = ???

jimthompson5910 (jim_thompson5910):

it might help to rewrite x+10 as x-(-10)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

wait would it be 10-k?

jimthompson5910 (jim_thompson5910):

\[\Large x + 10\\ \Large x - (-10)\\ \Large x - (\color{red}{-10})\\ \Large x - \color{red}{k} \] So k = -10

OpenStudy (anonymous):

but there isnt a -10 as an answer

jimthompson5910 (jim_thompson5910):

oh I went the wrong direction. It's to the right, so x gets replaced with x-6

OpenStudy (anonymous):

so it would be x+2?

jimthompson5910 (jim_thompson5910):

f(x) = 2^(x+4) f(x-6) = 2^(x-6+4) f(x-6) = 2^(x-2) so k = 2

jimthompson5910 (jim_thompson5910):

x-2 actually

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i got it right :D

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

np